OpenStudy (anonymous):

Angle A & Angle B are complimentary angles. If Angle A is 6 more than 3 tims Angle B what are the measures of the angles?

OpenStudy (anonymous):

A+B=90 A = 3B + 6 solve this system to get the measures of angle A and angle B.

OpenStudy (anonymous):

try substitution: A + B = 90 (3B+6) + B = 90 can you take it from here?

OpenStudy (anonymous):

not sure let me see

OpenStudy (anonymous):

where do you get 90 from